### v2

Abstract
Cyber-physical systems (CPSs) are complex systems that integrate physical, computational, and communication subsystems. The heterogeneous nature of these systems makes their safety assurance challenging. In this paper, we propose a novel automated approach for guardrailing cyber-physical systems using property-based tests (PBTs) generated by Large Language Models (LLMs). Our approach employs an LLM to extract properties from the code and documentation of CPSs. Next, we use the LLM to generate PBTs that verify the extracted properties on the CPS. The generated PBTs have two uses. First, they are used to test the CPS before it is deployed, i.e., at design time. Secondly, these PBTs can be used after deployment, i.e., at run time, to monitor the behavior of the system and guardrail it against unsafe states. We implement our approach in ChekProp and conduct preliminary experiments to evaluate the generated PBTs in terms of their relevance (how well they match manually crafted properties), executability (how many run with minimal manual modification), and effectiveness (coverage of the input space partitions). The results of our experiments and evaluation demonstrate a promising path forward for creating guardrails for CPSs using LLM-generated property-based tests.
Keywords:
Property-based Testing LLM4SE Cyber-Physical System Safety
1 Introduction
Cyber-physical systems (CPS) are integrated hardware-software systems where computation and physical processes are deeply intertwined. Ensuring safety [7] in these systems, in particular for the safety-critical ones is of high importance, as failures can have critical consequences. One of the key strategies in safety assurance is to capture the system’s properties describing what the system should and should not do under different conditions. Encoding the requirements as formal or semi-formal properties, can enable creating safety and security guardrails for system behavior. Formulating properties usually starts from the system requirements typically written in natural language. Therefore, large language models (LLMs) with their strong potential can be leveraged to extract properties from existing documentation and software code. These properties can be used to drive subsequent automated testing and verification activities, such as property-based testing [5, 3]. Property-based tests (PBT) are software tests that check a given property regarding the expected behavior holds for various input scenarios.
In this paper, we propose a novel automated and scalable approach for guardrailing CPS using PBT generated by LLMs. Our approach benefits from two major established facts in software engineering: 1) the cyber side of CPSs is essentially a software program amenable to existing automated program analysis tools; and 2) advanced LLMs are strong in analyzing programs and extracting their expected properties [20]. Based on these two observation, our proposed approach, called ChekProp, uses LLMs to generate property-based tests for CPSs before their deployment, i.e., at design time. These PBTs can also then be utilized after deployment, i.e., at run time, to detect unsafe behavior of the CPS. ChekProp uses the source code, documentation, and unit tests of the target CPS to extract properties regarding its expected behavior. ChekProp also generates PBTs that verify that the extracted properties hold for the CPS. We implement a prototype of ChekProp and make it publicly available to the community [4].
We evaluate the relevance of properties extracted by ChekProp for nine programs: two widely studied CPSs and seven Raspberry Pi programs. ChekProp extracts 25 properties on these nine programs. Our results show that the properties extracted by ChekProp are similar to those carefully created with manual effort, with a recall of 94% and a precision of 72%. The high precision and recall of ChekProp shows that it can be a reliable tool for automating the manual effort dedicated to property extraction for CPSs. Moreover, we study the quality of PBTs generated by ChekProp. We find that 47% of the PBTs generated by ChekProp become executable with minor modifications and 85% of the PBTs effectively cover various parts of the input space partitions. This suggests that ChekProp generates PBTs that successfully verify the CPS compliance with the extracted properties.
In summary, our main contributions are the following.
  • We propose a novel automated approach for generating property-based tests for CPSs using LLMs.
  • We implement a prototype of our proposed approach in ChekProp and make it accessible to the community in our open-source repository [4].
  • We report the results of our preliminary experiments on the relevance and quality of the PBTs generated by ChekProp in practice.
2 Background on Property-based Testing
Property-based testing was first introduced in QuickCheck [3]. Given a function under test $f$, the input space of this function $X$, and a property $P$ that checks the behavior of $f$ on a given input, a property-based test validates that $\forall x\in X:P(x,f)$. The property $P$ can be seen as a function that takes an input $x$ and the function $f$ and outputs true or false. The output of $P$ determines if $f$ behaves according to predefined requirements on $x$. In practice, a property-based test (PBT) consists of three components: 1) an input generator gen(), which returns different inputs, like $x$, from the input space $X$; 2) a test body that collects relevant data regarding the behavior of $f$; and 3) a test assertion that uses the data collected by the test body to assert that the property $P$ about $f$ is true for a given input $x$.
Listing 1: An example of a Python property-based test that checks the pow method returns a positive number as the square of an integer. This PBT uses the hypothesis library to generate inputs for the test.
1 The property-based unit test for the pow method:
2@given(st.integers())
3def test_pow2_positive_output(x):
4 square = pow(x, 2)
5 assert square >= 0
6
7 The example-based unit test for the pow method:
8def test_pow2_on_negative_input()
9 square = pow(-3, 2)
10 assert square == 9
Take 1 as an example. The PBT in this example (lines #1-5) tests the Python pow method. The main goal of this test is to check that the pow method returns a positive number when it powers an integer by 2. The PBT employs the hypothesis library [11] for input generation (line #2). The given decorator at line #2 uses the st.integers() strategy to generate various random integers and consider them as input $x$ at line #3. The hypothesis library provides the given decorator, the st.integers() strategy, and many other tools to facilitate property-based testing in Python. After the input is generated with the help of hypothesis library, the PBT in 1 collects the output of pow and saves it in square at line #4. Finally, it checks the property that the output of pow is positive at line #5. This property-based test delivers exactly what we need: a test that checks that the pow method returns a positive number for various integers, when they are powered by 2.
To better understand property-based testing, we can compare it with the commonly used example-based unit tests (EBTs), which test program with fixed arguments [18]. An EBT checks if the function $f$ under test works correctly for a single input $x$. For this, the EBT usually inspects that the output of $f$ for $x$ is exactly the same as the correct output $o$ determined by an oracle. In contrast, a PBT checks that $f$ behaves according to expectations for various inputs from the input space $X$.
1 shows an example of EBT for the Python pow method as well. The main goal of this test is also checking that the pow method returns a positive number when it powers an integer by 2. The EBT (lines #7-10) tests the pow method by giving it a negative integer, namely -3, as base, and 2 as power. Then, it checks that the output of pow is exactly 9 (line #10). If this test passes, it only shows that pow method returns a positive number as the square of -3 (as an example of an integer). The EBT is testing pow only for one single input, and its result might not be generalizable. Also, this type of testing requires an oracle that states the expected output, which is 9 in this case.
PBTs are suitable for safety checking the behavior of cyber-physical systems at runtime for two reasons. First, in PBTs, we do not need to know the exact expected behavior of the CPS under test. PBTs only check that the behavior of the CPS meets certain requirements, which can indicate the safety of its behavior. Secondly, PBTs check the behavior of the CPS on any input from the input space. This enables PBTs to ensure the safety of the CPS under unforeseen situations at runtime. We need to utilize the test body and test assertion of the PBT to monitor and assure the safety of CPS behavior at runtime. Based on these observations, in this paper, we use LLMs to generate PBTs for cyber-physical systems.
3 Proposed Approach
We envisage a novel approach for gaurdrailing cyber-physical systems with LLM-generated PBTs. Figure 1 illustrates an overview of our proposed approach. This approach consists of two main phases: the PBT generation phase, and the property-based monitoring phase.
The PBT generation phase occurs at design time, before the system is deployed in real world. In this phase, we employ LLMs to generate PBTs that guardrail cyber-physical systems against running in unwanted/unsafe states. We implement our tool ChekProp to carry out the proposed PBT generation, given the documents, code, and unit tests of the cyber-physical system. After the PBTs are generated, we enter the property-based monitoring phase of our approach. This phase occurs at run time, when the system is deployed in real world. In this phase, the running cyber-physical system is constantly checked against the generated PBTs. Once a violation of a PBT is detected, a warning is raised and the proper safety measures should be taken.
Figure 1: Overview of the proposed two-phase approach. ChekProp particularly focuses on the PBT generation phase. In the PBT improvement loop step, ChekProp aims to improve the generated PBTs by sending the LLM an improvement prompt, consisting of the failed PBTs and the error messages collected for them. Errors of various types are considered, including syntax errors, compilation errors, exceptions thrown during test executions, and assertion failures.
In the following, we explain the proposed PBT generation method, which is implemented in ChekProp. ChekProp takes the natural language documents of the CPS, its source code, and unit tests and prompts an LLM to generate PBTs. Next, it analyzes existing PBTs to detect issues and iteratively prompts the LLM to improve the PBTs. Once the PBTs pass the analysis, they are considered as verified PBTs that can be used as guardrails for the CPS. We now discuss each of these ChekProp components in more detail.
3.1 Inputs of ChekProp
The input to ChekProp comprises natural language documents that specify the CPS and its expected behavior, the CPS source code in Python, and unit tests for this source code.
The natural language documents of the CPS describe the expected behavior of the system and the constraints that should be observed in the run time. Take Figure 2 as an example of a natural language document that describes a Pneumatic Control System (PCS) [12]. The description first defines the main elements involved in the system, namely, the horizontal and vertical cylinders and their corresponding sensors and controllers. Next, it explains the expected behavior of the system and the expected order of cylinder movements. Finally, it presents the constraints that should be met during the movements.
Given the natural language documents of the system, the CPS is implemented to follow the described requirements. Moreover, a set of unit tests is created to test the CPS implement for specific points in the input space. Note that implementing the CPS and creating unit tests for it can also be fully automated using state-of-the-art LLM-based code generation [8] and test generation [21] techniques. However, ChekProp focuses on PBT generation and assumes that the CPS implementation is provided in Python, along with at least one unit test that demonstrates how the test body should interact with different methods of the program.
Figure 2: The natural language document that describes a Pneumatic Control System (PCS). We take the original design of PCS from [12] and adapt it to make it suitable for property-based testing.
3.2 Initial PBT Generation
ChekProp starts PBT generation by synthesizing an initial prompt. This prompt is used to invoke the LLM for generating an initial batch of PBTs. The initial prompt consists of four main sections and follows the structure presented in Figure 3.
As illustrated in the Figure 3 example, the first section presents the description of the CPS in natural language. LLMs are highly effective in understanding natural language specifications of software and translating those specifications to actual code [8]. Therefore, we provide this section of the prompt to help the LLM better understand the system constraints that should be later translated into PBTs.
The second part of the initial prompt contains the Python code for the CPS. In Figure 3, the second section presents a part of the code for pneumatic control system, namely, the Cylinder class (line #14). Including the system code in the prompt is essential for the LLM to recognize how the system should be called in tests.
The third part of the initial prompt also provides at least one example unit test for the system. The third part of Figure 3 shows an example of a unit test that calls the system controller. This unit test employs an instance of the MockSystem class (line #58) to mock the physical part of the pneumatic control system and obtain a simple interface to its controller. It also illustrates how the states of the system should be collected during execution and checked later (lines #59-61).
Note that, as explained in section 2, there is significant difference between unit tests and property based tests. The unit test only checks the behavior of the program for a specific input. For example, in the unit test (Section 3) of Figure 3, specific total_time, cylinder_interval, etc. are used. Also, in unit testing we usually check that the output is exactly what is expected according to an oracle [19]. In contrast, property based tests check that a more general condition is meet by the program behavior over a wide range of inputs.
The fourth and final part of the Figure 3 instructs the LLM to generate the desired PBTs. At the end of the initial PBT generation step, ChekProp obtains a set of initial PBTs. ChekProp runs these PBTs and collects their results using an analyzer unit. If a group of generated PBTs fails, ChekProp collects their failure message and enters its PBT improvement loop step as described in subsection 3.3.
Figure 3: The structure of initial prompt used by ChekProp for generating an initial set of PBTs.
3.3 PBT Improvement Loop
In the PBT improvement loop step, ChekProp aims to improve the suite of generated PBTs. For this, ChekProp sends the LLM an improvement prompt, consisting of the failed PBTs and the error messages collected for them. Errors of various types are considered, including syntax errors, compilation errors, exceptions thrown during test executions, and assertion failures. The improvement prompts are sent to the LLM in continuation of the initial prompt, which means that the LLM also has the CPS description, code, and unit test in context. Therefore, the LLM has all the information needed to improve the PBTs.
The PBT improvement loop component of ChekProp, iteratively sends improvement prompts to the LLM and employs the analyzer unit to run the improved PBTs and collect their results. If the improved PBTs still fail, ChekProp repeats this process until all PBTs are fixed or ChekProp reaches a predefined maximum number of improvement attempts.
3.4 Output of ChekProp
Listing 2: The PBT generated by ChekProp for the Pneumatic Control System presented in Figure 3  with three parts, the input generator, the test body, and the test assertions.
1from hypothesis import given, strategies as st
2from examples.gpiozero.apps.pcs.src.pcs import MockSystem
3
4
5# Test that the cylinder locations remain within the bounds (0 and 2) at all times.
6@given( # This is the input generator of the PBT, which uses the hypothesis library to run the test with different inputs.
7 total_time=st.floats(min_value=1.0, max_value=100.0),
8 cylinder_interval=st.floats(min_value=0.1, max_value=10.0),
9 controller_interval=st.floats(min_value=0.1, max_value=10.0),
10 mock_interval=st.floats(min_value=0.1, max_value=10.0),
11)
12def test_cylinder_location_in_bounds(total_time, cylinder_interval, controller_interval, mock_interval):
13 system = MockSystem(total_time, cylinder_interval, controller_interval, mock_interval) # Start of test body
14 collected_states = system.execute_scenario()
15
16 for state in collected_states: # End of test body
17 # The following two lines present the test assertions of this PBT
18 assert 0 <= state.cylinder_a_loc <= 2, f"Cylinder A out of bounds: {state.cylinder_a_loc}"
19 assert 0 <= state.cylinder_b_location <= 2, f"Cylinder B out of bounds: {state.cylinder_b_location}"
In the event of a successful PBT generation, ChekProp outputs the PBT as Python code. In particular, per our experiments, LLMs always use the hypothesis library [11] to write property based tests in Python. The tests generally use the same testing framework as the example unit test. For example, the unit test in Figure 3 is run with pytest. The PBTs generated by LLMs for this prompt can also be executed with pytest.
2 shows one of the PBTs that ChekProp generates for the pneumatic control system. As the comment above the test mentions, it checks that cylinders stay within the location bound. The property is checked by two assertions at lines #17-18, one assertion per cylinder. This property is checked for a range of different settings for cylinders and their controllers presented at lines #7-10. Note the main difference between this PBT and the unit test in Figure 3: the unit test checks that a specific output for a given input is exactly correct, while the PBT verifies that a general property holds for all inputs within a specified range. This makes PBTs more general and appropriate for checking that the system does not show unsafe behavior in unforeseen situations.
3.5 Implementation
ChekProp uses the gemini-2.0-flash-lite-preview-02-05 in its current version, but adopts a flexible design that allows easy switch to other LLMs. ChekProp also invokes the LLM with a sample size of one and a temperature of zero, which means that it receives only the top response per LLM invocation. In the current version of ChekProp, the PBT improvement loop is disabled, and we assess LLM’s ability to generate PBTs at the initial attempt.
3.6 Property-based Monitoring
While ChekProp is focused on PBT generation (phase 1 in Figure 1), property-based monitoring (phase 2 in Figure 1) is also an integral part of our guardrailing CPSs with PBTs. In the monitoring phase of our proposed approach, various components of the generated PBTs are used to collect relevant data and verify the properties at runtime. For example, the generated PBT in 2 tests that the cylinders stay in the correct location range for given inputs. This test shows that we should collect state values (line #16) and then assert that their cylinder_a_loc attribute is in the range $[0,2]$ (line #17) to ensure the property holds for cylinder A. In the monitoring phase, we can use the same data collection and property assertion techniques to check that the cylinders do not enter unsafe locations at runtime. This example demonstrates how the property extracted by ChekProp and the implementation of its generated PBTs are useful, relevant, and vital for guardrailing CPSs.
More generally, as explained in section 2 and as seen in 2, the PBTs generated in the first phase consist of three components: an input generator, a test body, and a test assertion. In the property-based monitoring phase, the PBT input generator is no longer needed, as the inputs are generated by the cyber part (in the form of control commands) and the physical system (via sensor values). The test body is replaced by the monitor that sits between the controller and the physical system collecting relevant data. The properties derived at design time are transformed into guards that are checked at runtime. The monitor verifies that these guards hold in the current state of the system. Based on the situation, if a violation is detected, the monitor can intercept and block the command being sent to the physical system. In this way, the PBTs generated by ChekProp serve as runtime guardrails for CPSs.
4 Experiments
4.1 Research Questions
We conduct preliminary experiments to answer the following research questions:
In this paper, we study the quality of ChekProp extracted properties and the quality of its generated property-based tests according to the following research questions:
  • RQ1 (Property relevance): Does the proposed approach extract relevant properties? We assess the quality of properties extracted by our approach on a dataset of Python programs for cyber-physical systems. Our dataset consists of two cyber-physical systems that are extensively studied in the literature, as well as seven Raspberry Pi programs. We compare the automatically extracted properties with manually crafted properties to judge their relevance.
  • RQ2 (PBT quality): What is the quality of ChekProp generated PBTs for real-world CPSs? We use ChekProp to generate PBTs for Python CPS programs in our dataset. We assess the quality of ChekProp generated PBTs from two aspects. First, we check if the generated PBTs can be executed with minimal manual modification (executability). Second, we evaluate the extent of various input space partitions on which the generated PBTs execute the program (effectiveness in terms of coverage of the input space partitions).
4.2 Dataset
Table 1: The cyber-physical system Python programs considered in our dataset.
ID Program Type Description
P1 Temperature Control
System (TCS)
Model-based
Example [12]
The TCS system controls the temperature of a room. The room has a window that blows cold or hot wind into the room. A sensor collects the current temperature and a HC unit is used to heat up or cool down the room environment. Finally, a controller collects the temperature data from the sensor and decides if the HC unit should be used.
P2 Pneumatic Control
System (PCS)
Model-based
Example [12]
The PCS regulates the movement of components of a mechanical system that consist of a horizontal and a vertical cylinder, their corresponding controllers and sensors. This mechanical system is designed to pick an object from the ground and move it to a new place.
P3 Laser Tripwire Raspberry
Pi Projects [15]
A prototype on a breadboard to detect whether a beam of light is hitting the light-dependent resistor (LDR).
P4 Line Following Robot Raspberry
Pi Projects [15]
A robot with two motors and two sensors. The robot uses its right and left sensors to follow a line and decide if should go turn right, left, or go forward.
P5 Ultrasonic Theremin Raspberry
Pi Projects [15]
A system that produces a sound with a volume level corresponding to the user distance from the device.
P6 Remote Buggy Raspberry
Pi Projects [15]
A robot that moves in four directions and a user remotely controls it.
P7 Quick Reaction Game Raspberry
Pi Projects [15]
A game in which a light turns on and off and two players compete to hit their button faster after the light is turned off.
P8 Presence Indicator Raspberry
Pi Projects [15]
A monitor made of a list of LEDs that shows the number of people present at a place.
P9 gpiozero InputDevice gpiozero
Library Class[17]
A core class in gpiozero library that handles connection to input devices on a Raspberry Pi board.
As the current version of ChekProp supports PBT generation for Python programs, we curate a dataset of Python CPS programs for our experiments. This dataset consists of nine programs presented in Table 1. These programs are taken from three main sources as follows.
First, we include the Python version of two CPSs that are widely studied in the model checking literature [12]: a temperature control system (TCS) and a pneumatic control system (PCS) (P1 and P2 in Table 1). TCS and PCS are presented with the IDs P1 and P2 in Table 1. Moradi et al. [12] use the model checking tool of Rebeca  [^1] Rebeca is an actor-based modeling language, in which actors are the units of concurrency and communicate via asynchronous message passing [16]., Afra [16], to detect potential attacks against these systems. For this, they define the correctness properties for each system and assess if Afra can find counterexamples for these properties on models augmented with malicious behavior. The manually defined properties in [12] set a ground-truth with which we can compare the properties automatically extracted by ChekProp.
We carefully implement TCS and PCS in Python to make them amenable to PBT generation by ChekProp. 3 shows a summary of our implementation of the TCS. There is a class for each component (i.e. Rebeca actor) of the system, namely, TempSensor for the sensor, HCUnit for the HC unit, and Controller for the controller. Each component runs on a separate thread and updates its status periodically. For example, the sensor fetches current temperature every sensor_interval seconds (see line #17). We also provide a MockRoom class that simulates the room environment and enables us to execute the system with different configurations, such as initial_temp and sensor_interval (see lines #50 and #58). This implementation is suitable for testing the temperature control system.
The second set of programs in our dataset are the six programs (P3-P8 in Table 1) taken from open-source Raspberry Pi projects [15] that use the gpiozero library [17]. gpiozero is a Python library that real-world cyber-physical systems employ to connect to Raspberry Pi boards. We take the source of these programs from the official Raspberry Pi website [15] and manually add unit tests for them. These unit tests can be used in ChekProp prompts (see subsection 3.1). The six programs in our dataset that use gpiozero enable us to evaluate ChekProp applicability on projects that adopt the widely used Raspberry Pi boards.
Listing 3: Implementation of the temperature control system in Python.
1class Environment:
2 def __init__(self, initial_temp: int = None):
3 self.temp = initial_temp if initial_temp is not None else random.randint(20, 24)
4
5 def fetch_temp(self):
6 return self.temp
7 ....
8
9class TempSensor:
10 def __init__(self, env: Environment):
11 self.env = env
12 self.temp = self.env.fetch_temp()
13
14 def start_temp_collection(self, total_time: float, sensor_interval: float):
15 for i in range(math.floor(total_time / sensor_interval)):
16 self.temp = self.env.fetch_temp()
17 sleep(sensor_interval)
18
19class PWMOutputDevice:
20 .....
21
22class HCUnit:
23 def __init__(self):
24 self.cooler = PWMOutputDevice()
25 self.heater = PWMOutputDevice()
26
27 def activate_cooler(self):
28 self.cooler.on()
29 self.heater.off()
30 ....
31
32class Controller:
33 def __init__(self, temp_sensor: TempSensor, hc_unit: HCUnit):
34 self.temp_sensor = temp_sensor
35 self.hc_unit = hc_unit
36
37 def control(self, total_time: float, control_interval: float):
38 for i in range(math.floor(total_time / control_interval)):
39 temperature = self.temp_sensor.temp
40 if 21 <= temperature <= 23:
41 self.hc_unit.deactivate()
42 ...
43
44class SystemState:
45 def __init__(self, temp, cooler_state, heater_state, outside_air_temp):
46 self.temp = temp
47 ...
48
49class MockRoom:
50 def __init__(self, total_time: float, sensor_interval: float, control_interval: float, initial_temp: int = None):
51 self.env = Environment(initial_temp=initial_temp)
52 self.total_time = total_time
53 self.sensor_interval = sensor_interval
54 self.control_interval = control_interval
55 self.temp_sensor = TempSensor(self.env)
56 ...
57
58 def execute_scenario(self):
59
60 sensor_thread = threading.Thread(target=self.temp_sensor.start_temp_collection,
61 args=(self.total_time, self.sensor_interval))
62 sensor_thread.start()
63 ...
64
65 collected_states = []
66 for i in range(self.total_time):
67 ...
68 outside_air_temp = self.env.get_outside_air_temp()
69 collected_states.append(SystemState(cur_temp, ...))
70 self.env.set_temp(cur_temp + outside_air_temp + heater_value - cooler_value)
71 sleep(1)
72
73 sensor_thread.join()
74 control_thread.join()
75
76 return collected_states
Listing 4: The InputDevice class in gpiozero.
1class InputDevice(GPIODevice):
2 """
3 Represents a generic GPIO input device.
4
5 This class extends :class:‘GPIODevice‘ to add facilities common to GPIO
6 input devices. The constructor adds the optional *pull_up* parameter to
7 specify how the pin should be pulled by the internal resistors. The
8 :attr:‘is_active‘ property is adjusted accordingly so that :data:‘True‘
9 still means active regardless of the *pull_up* setting.
10
11 :type pin: int or str
12 :param pin:
13 The GPIO pin that the device is connected to. See :ref:‘pin-numbering‘
14 for valid pin numbers. If this is :data:‘None‘ a :exc:‘GPIODeviceError‘
15 will be raised.
16
17 :type pull_up: bool or None
18 :param pull_up:
19 If :data:‘True‘, the pin will be pulled high with an internal resistor.
20 If :data:‘False‘ (the default), the pin will be pulled low. If
21 :data:‘None‘, the pin will be floating. As gpiozero cannot
22 automatically guess the active state when not pulling the pin, the
23 *active_state* parameter must be passed.
24
25 :type active_state: bool or None
26 :param active_state:
27 If :data:‘True‘, when the hardware pin state is ‘‘HIGH‘‘, the software
28 pin is ‘‘HIGH‘‘. If :data:‘False‘, the input polarity is reversed: when
29 the hardware pin state is ‘‘HIGH‘‘, the software pin state is ‘‘LOW‘‘.
30 Use this parameter to set the active state of the underlying pin when
31 configuring it as not pulled (when *pull_up* is :data:‘None‘). When
32 *pull_up* is :data:‘True‘ or :data:‘False‘, the active state is
33 automatically set to the proper value.
34
35 :type pin_factory: Factory or None
36 :param pin_factory:
37 See :doc:‘api_pins‘ for more information (this is an advanced feature
38 which most users can ignore).
39 """
40 def __init__(self, pin=None, *, pull_up=False, active_state=None,
41 pin_factory=None):
42 super().__init__(pin, pin_factory=pin_factory)
43 try:
44 self.pin.function = ’input’
45 pull = {None: ’floating’, True: ’up’, False: ’down’}[pull_up]
46 if self.pin.pull != pull:
47 self.pin.pull = pull
48 except:
49 self.close()
50 raise
51
52 if pull_up is None:
53 if active_state is None:
54 raise PinInvalidState(
55 f’Pin {self.pin.info.name} is defined as floating, but
56 f’"active_state" is not defined’)
57 self._active_state = bool(active_state)
58 else:
59 if active_state is not None:
60 raise PinInvalidState(
61 f’Pin {self.pin.info.name} is not floating, but
62 f’"active_state" is not None’)
63 self._active_state = False if pull_up else True
64 self._inactive_state = not self._active_state
65
66 @property
67 def pull_up(self):
68 """
69 If :data:‘True‘, the device uses a pull-up resistor to set the GPIO pin
70 "high" by default.
71 """
72 pull = self.pin.pull
73 if pull == ’floating’:
74 return None
75 else:
76 return pull == ’up’
77 ...
The last program considered in our dataset (P9 in Table 1) is InputDevice, a core class from the gpiozero library. This class provides an interface for Python programs to interact with Raspberry Pi input devices, such as barometers, temperature sensors, etc. The InputDevice class is a complex class from gpiozero that connects to various components in this library. Testing this class requires a detailed understanding of the inner workings of the library and how it represents and handles the physical environment. For example, to instantiate an object of the InputDevice class, we have to pass the number of a pin to its constructor method. The type of this pin should be “input”, while some of the pins on a Raspberry Pi board are reserved only for “output”. A correct test should use a pin of the correct type to instantiate an InputDevice. Another example of such details in gpiozero is how pins are activated. To activate a pin on a Raspberry Pi board, its voltage should go high, which can happen by calling the pin.drive_high() method. Writing a test for a CPS program requires an accurate understanding of how interacting with the program, such as calling pin.drive_high(), impacts the physical status of the system, such as increasing the voltage on the Raspberry Pi board. To evaluate whether ChekProp generated PBTs capture such details about CPS programs, we include the InputDevice class in our dataset.
4 presents the InputDevice class. This class, similar to other gpiozero classes, has well-written documentation (lines #2-39). We consider this documentation as the natural language description in ChekProp prompts (see subsection 3.1). Moreover, gpiozero has extensive unit tests for its classes, which we use them to produce our initial prompt . This confirms that gpiozero classes have the essential components for applying ChekProp: the natural language description, the source code, and the unit test (see subsection 3.1).
Overall, our dataset contains a combination of CPSs studied in research literature and CPS programs that employ widely used libraries. This dataset helps us to assess the relevance of ChekProp extracted properties and the quality of its generated PBTs.
4.3 RQ1: Property Relevance
4.3.1 I. Methodology:
To assess the quality of extracted properties, we run our approach on the nine programs in our dataset and analyze the relevance of the extracted properties. In this experiment, we abstract away the implementation details of the generated PBTs; instead, we focus on how well the properties considered in the PBTs validate the logic of the program under test. For this purpose, we compare properties extracted by our approach with manually crafted properties that we consider as ground-truth. In particular, we evaluate whether the logic checked by ground-truth properties is also validated by ChekProp extracted properties and vice versa. As explained in subsection 4.2, for programs P1 and P2, the ground-truth properties are already stated by Moradi et al. [12]. For the remaining programs (P3-P9), we manually define the ground-truth properties.
Take the temperature control system (TCS) as an example. Moradi et al. outline three properties for TCS as follows:
  1. 1.
    If the room is warm (temp > 23), the HC unit should not be heating the room.
  2. 2.
    If the room is cold (temp < 21), the HC unit should not be cooling the room.
  3. 3.
    The temperature should never be too low (temp < 20) or too high (temp > 24).
We first apply our proposed approach on our Python implementation of TCS to generate PBTs. Next, we compare the extracted properties that are tested in these PBTs with the three ground-truth properties in [12]. If the ChekProp properties correspond with the three ground-truth properties, we conclude that the proposed approach is able to extract useful properties.
4.3.2 II. Results:
LABEL:tab:rq1_res shows the results of this experiment. In total, the table contains 26 properties. We split these properties into four groups: Group1 consists of 15 properties that are present among ground-truth and ChekProp extracted properties in the exact same form ( Pr3, Pr5, Pr7, Pr9, Pr10, Pr11, Pr12, Pr14, Pr15, Pr17, Pr20, Pr21, Pr23, Pr24, and Pr25); Group2 consists of 3 properties that are present among ground-truth and ChekProp extracted properties in equivalent but slightly different forms (Pr1, Pr2, and Pr6); Group3 consists of 7 properties that are only among the ChekProp extracted properties (Pr4, Pr8, Pr13, Pr16, Pr19, Pr22, and Pr26); and Group4 consists of 1 property that is only among ground-truth properties (Pr18).
In total, the ground-truth contains 19 properties (Group1+Group2+Group4) and ChekProp extracts 25 properties (Group1+Group2+Group3). Among all properties, 18 are common between ground-truth and ChekProp, either in the exact same form (Group1) or with distinct different formulations (Group2). These properties are relevant, since they are present among the manually crafted properties. Therefore, the recall of ChekProp is 94% (18/19), which indicates that our approach can fully replace the manual effort required for extracting most of the properties from CPSs. The precision of ChekProp is 72% (18/25), suggesting that the properties extracted by ChekProp often represent what humans expect from the CPS under test. The high precision and recall of ChekProp make it a reliable tool for automating the manual effort dedicated to property extraction for CPSs.
Table 2: Comparison between properties automatically extracted by our proposed approach and ground-truth properties that are manually crafted.
ID Program Ground-truth Property Corresponding ChekProp Property
Pr1 TCS If the room is warm (temp > 23), the HC unit should not be heating the room. Heater should be activated if the temperature drops below 21°C.
Pr2 TCS If the room is cold (temp < 21), the HC unit should not be cooling the room. Cooler should be activated if the temperature exceeds 23°C.
Pr3 TCS The temperature should never be too low (temp < 20) or too high (temp > 24). Same as ground-truth.
Pr4 TCS Neglected. Neither heater nor cooler is active when temperature is in the target range (21-23°C).
Pr5 PCS The horizontal cylinder should not move when the vertical cylinder is down. Same as ground-truth.
Pr6 PCS The horizontal and vertical cylinders should not move simultaneously. The movement should follow a specific order.
Pr7 PCS The cylinders location should always be between 0 and 2. Same as ground-truth.
Pr8 PCS Neglected. The cylinder movement speed should not exceed 1.
Pr9 Laser Tripwire "INTRUDER" is printed if an only if there is no light. Same as ground-truth.
Pr10 Line Following Robot When left sensor is on, left motor goes backwards and right motor goes forward. Same as ground-truth.
Pr11 Line Following Robot When right sensor is on, right motor goes backwards and left motor goes forward. Same as ground-truth.
Pr12 Line Following Robot When both sensors are off, both motors go forward. Same as ground-truth.
Pr13 Line Following Robot Correctly not included. When both sensors are on, at least one motor is moving.This is an irrelevant property that never occurs in practice.
Pr14 Ultrasonic
Thermin
Volume level never gets too high or too low. Same as ground-truth.
Pr15 Ultrasonic
Thermin
Volume increases when user is getting closer to the sensor. Same as ground-truth.
Pr16 Ultrasonic
Thermin
Correctly not included. The buzzer has a correct range of volume.This is an irrelevant property that mostly checks the inner classes of the gpiozero library, not the application.
Pr17 Remote Buggy Pressing each button on the controller, activates the corresponding motor. Same as ground-truth.
Pr18 Quick Reaction Game Check that the light turns on at some point and turns off afterwards. Not extracted.
Pr19 Quick Reaction Game Correctly not included. When the valid_pressed method is called, the name of the winner is printed.This is not a useful property, as it only tests very detailed implementation details.
Pr20 Presence
Indicator
The LEDs show the number of present people divided by 10. Same as ground-truth.
Pr21 Presence
Indicator
When the number of present people is above 10, the number “1” is shown. Same as ground-truth.
Pr22 Presence
Indicator
Correctly not included. When the PresneceIndicator object is closed, the LEDs are off.This is not a useful property, as it only tests very detailed implementation details.
Pr23 gpiozero
InputDevice
The pull_up parameter of the input device affects the pull state of the pin. Same as ground-truth.
Pr24 gpiozero
InputDevice
With a none pull_up, the active_state should be set. Same as ground-truth.
Pr25 gpiozero
InputDevice
When pull_up is set, the is_active parameter has a reverse effect between the pull_up value of the device and the activation of the pin. Same as ground-truth.
Pr26 gpiozero
InputDevice
Neglected. When the input device is closed, the pin is not in use anymore.
In LABEL:tab:rq1_res, we see that in three cases (Pr4, Pr8 and Pr26) ChekProp extracts a property that is relevant and useful, but neglected in manually crafted properties. For example, our approach extracts Pr4 for PCS which notes that neither heater or cooler should be active when the room temperature is between 21°C and 23°C. This indicates that not only can our automated approach replace manual property design work, but it can also even improve the manually crafted properties.
The three relevant properties neglected in ground-truth together with the 18 properties common between ground-truth and ChekProp make up the set of our 21 relevant properties.
As presented in LABEL:tab:rq1_res, four of the properties extracted by ChekProp (Pr13, Pr16, Pr19, and Pr22) are not useful. These properties either check a state that does not occur in real-wrold (Pr13) or validate highly detailed implementation nuances. This observation shows that a manual check on properties automatically extracted by ChekProp is needed to ensure that no useful property is considered for testing.
Finally, there is only one ground-truth property (Pr18) that does not correspond to any of the properties extracted by ChekProp. Pr18 is a property for the quick reaction game and indicates the order of changes in the light status, it should be first turned on at some point and then turned off at some point. With a careful manual analysis, we understand that the code we provide to the LLM for the quick reaction game lacks the documentation regarding this point. This suggests the importance natural language description as of one core components in ChekProp prompts (see subsection 3.1).
{mdframed}
Answer to RQ1: Does the proposed approach extract relevant properties?
We compare the manually crafted relevant properties with ChekProp extracted properties for nine programs in our dataset. This comparison shows that 94% (18/19) of the ground-truth properties are also automatically extracted by ChekProp. Moreover, ChekProp extracts three additional relevant properties that are neglected in manually crafted properties. This indicates that ChekProp is a reliable tool for automating the tedious and complicated task of defining CPS properties..
4.4 RQ2: PBT Quality
4.4.1 I. Methodology:
For evaluating the quality of PBTs generated by ChekProp, we examine the PBTs that test the 21 relevant properties according to our analysis in the RQ1 experiment (see subsection 4.3). As explained in subsection 4.1, we assess the applicability of our approach from two aspects: executability and effectiveness.
We consider a PBT executable if and only if it is correct both syntactically (i.e., successfully compiles) and semantically (i.e. passes). To investigate the executability of a PBT, we check to what extent the PBT should be manually modified to reach syntax and semantic correctness. A lower level of manual modification indicates higher executability and vice versa. We perform a manual analysis to find the level of executability of PBTs. Based on this analysis, we assign the PBTs generated for each program to one of the following executability levels: “HIGH”, “MED”, and “LOW”. A “HIGH” executability level means that the analyzer has to spend less than one minute manually fixing the PBT to ensure it runs and passes successfully. “LOW” means that more than three minutes of manual work is needed, and “MED” means that between one and three minutes is required. In this investigation, for the PBTs generated per each program, we also take note of the main challenges that require manual modifications. The results reveal potential opportunities for future improvement in ChekProp.
To assess the effectiveness of generated PBTs, we study if it checks the property over representatives of all or most partitions of the input space. As explained in section 2, one of the main components of a PBT is an input generator that produces various inputs from the input space. In this experiment, we determine to what extent the input generators of generated PBTs produce inputs from all partitions of the input space. The more partitions of input space considered by a PBT, the more effective the PBT is. We assess the effectiveness of PBTs generated for each program through a manual analysis and assign them to one of the three effective groups “HIGH”, “MED”, and “LOW”.
4.4.2 II. Results:
LABEL:tab:rq2_res summarizes the result of our experiment on ChekProp applicability. The “Property_ID” and “Program” columns indicate the property and the program that the PBT is testing. Note that the ID of the property is taken from LABEL:tab:rq1_res, which lists the properties extracted by ChekProp. The “Executability” column shows the result of our assessment of the executability of generated PBTs in terms of their syntactical and semantical correctness. The fourth column presents the main executability challenge of generated PBTs that should be addressed manually. Finally, the last column presents the level of effectiveness of PBTs generated for each program.
Table 3: The quality of PBTs generated by ChekProp for the 21 relevant propertie. A "HIGH", "MED", or "LOW" level in the "Executability" column indicates the PBT can be successfully executed with less than one minute, between one to three minutes, or more than three minutes of manual effort for modification, respectively. The "Effectiveness" column indicates the level of input space partitions covered by the PBT.
Property_ID Program Executability Main Executability Challenge Effectiveness
Pr1 TCS HIGH No major executability issues detected. HIGH
Pr2 TCS HIGH No major executability issues detected. HIGH
Pr3 TCS HIGH No major executability issues detected. HIGH
Pr4 TCS HIGH No major executability issues detected. HIGH
Pr5 PCS HIGH No major executability issues detected. HIGH
Pr6 PCS HIGH No major executability issues detected. HIGH
Pr7 PCS HIGH No major executability issues detected. HIGH
Pr8 PCS HIGH No major executability issues detected. HIGH
Pr9 Laser Tripwire HIGH No major executability issues detected. MED
Pr10 Line Following Robot LOW Wrong parameter passed to the pin.drive_up() method. HIGH
Pr11 Line Following Robot LOW Wrong parameter passed to the pin.drive_up() method. HIGH
Pr12 Line Following Robot LOW Wrong parameter passed to the pin.drive_up() method. HIGH
Pr14 Ultrasonic Thermin MED The used mocking method does not work for parameterized tests. HIGH
Pr15 Ultrasonic Thermin MED The used mocking method does not work for parameterized tests. HIGH
Pr17 Remote Buggy HIGH Exceeds timeout as tested on too many inputs. HIGH
Pr20 Presence
Indicator
MED The used mocking method does not work for parameterized tests. MED
Pr21 Presence
Indicator
MED The used mocking method does not work for parameterized tests. MED
Pr23 gpiozero
InputDevice
MED The used mocking method does not work for parameterized tests. HIGH
Pr24 gpiozero
InputDevice
MED The used mocking method does not work for parameterized tests. HIGH
Pr25 gpiozero
InputDevice
LOW The used mocking method does not work for parameterized tests.
The pin state is set with an incorrect use of the interface.
HIGH
Pr26 gpiozero
InputDevice
MED The used mocking method does not work for parameterized tests. HIGH
For 47% (10/21) of the relevant properties (Pr1, Pr2, Pr3, Pr4, Pr5, Pr6, Pr7, Pr8, Pr9, and Pr17), the generated PBTs are executable without major changes that require less than one minute of manual work. In fact, for none of these PBTs, except for the Pr17 PBT, no major executability issues are detected. These PBTs successfully execute and pass with little to none manual modification. Also, for Pr17 PBT, the problem is that the generated PBT runs the program for too many inputs, leading to a timeout. A developer who knows the logic of Pr17 property of the Remote Buggy program can fix the generated PBT by only modifying the number of random inputs that should be considered. Given the complexity of predicting the time needed for running a test on a CPS, this case shows the importance and positive impact of keeping a human in the loop of LLM-based PBT generation. In sum, our analysis of the executability of PBTs generated for Pr1, Pr2, Pr3, Pr5, Pr6, Pr7, Pr9, and Pr17 shows that for a remarkable number of relevant CPS properties ChekProp generates a PBT that is executable with minor manual modifications.
For seven of the properties (Pr14, Pr15, Pr20, Pr21, Pr23, Pr24, and Pr26), the only main challenge to executability of generated PBTs occurs in their mocking of the CPS. This challenge occurs because the mocking method employed has a conflict with property-based testing of the CPS under test. In particular, every time the test is executed for a specific input, all the pins used in the mock of the CPS should be initialized from scratch. However, the mocking method used in these PBTs only initializes the mock object once for all inputs considered in the PBT. This leads to a semantic problem with the logic of the CPS under test, as well as a syntactic error in using the hypothesis library. Consequently, these seven PBTs require a medium level of manual modification to become executable.
We notice that mocking CPS is both tricky and essential for testing. As CPSs are supposed to run in a physical environment, we need to mock how the environment affects CPS programs. This can require a detailed understanding of the relationship between various components of given CPS programs. Previous work shows that such domain knowledge can be effectively provided to LLMs by in-context learning, i.e., adding relevant examples to the prompt [6]. Based on this observation, we suggest that practitioners use a few-shot prompt with mocking examples to generate PBTs for cyber-physical systems with LLMs.
Finally, one of the main issues with generated PBTs for four properties (Pr10, Pr11, Pr12, and Pr25) is how they use gpiozero. For example, the PBT generated for Pr25 uses output pins of the Raspberry Pi board to initialize their object of the InputDevice. With further analysis, we realize that fixing the issues in this PBT depends on a deep understanding of multiple gpiozero classes. However, the current version of ChekProp only includes the documentation of the InputDevice class in the prompt. This documentation is taken from the comments presented in 4. To fix the issue with this PBT, the LLM also needs to have the documentation for other classes, such as PiGPIOFactory. We conclude that a strong LLM-based PBT generation tool for CPS programs requires augmenting prompts with all relevant information from the program documents.
Our analysis of the effectiveness of the generated PBTs shows that the PBTs generated for 85% (18/21) of the properties are highly effective; these PBTs test the property on most partitions of the input space. With a more detailed look, we observe that the generated tests tend to be more effective when the tests employ a straightforward and flexible mock of the CPS components. For example, as shown in 3, our Python implementation of TCS provides a MockRoom class. This class enables a tester to run the program with many different inputs only by changing a few parameters regarding the starting temperature of the room and the timing of updating various components. Using this mock class, the generated PBTs for TCS properties run the program with different configurations that represent all partitions of the input space. This experiment also reaffirms the importance of using flexible mocks with a straightforward API for testing CPSs.
{mdframed}
Answer to RQ2: What is the quality of ChekProp generated PBTs for real-world CPSs?
We assess the applicability of PBTs generated by ChekProp 21 relevant properties in our dataset from two aspects: executability and effectiveness. Our results reveal that a remarkable number of the generated PBTs are highly executable (47%) and highly effective (85%), which indicates the applicability of ChekProp. Our analysis also leads to two major suggestions for generating high-quality PBTs for CPSs with LLMs. First, the LLM should be aided in employing straightforward and flexible mocking by providing well-designed few-shot examples in the prompt. Secondly, it is important to include the relevant documentation from all parts of the CPS in the prompt. These two techniques make our proposed approach even more robust and practical.
5 Related Work
The application of LLMs to testing CPS is at early stages and many of the efforts have been focused on scenario generation for autonomous driving and robotics. For instance, OmniTester [10] uses an LLM (GPT-4) to generate diverse driving scenarios from natural language descriptions and proposes test road layouts and events. They also incorporates retrieval-augmented generation (RAG) and iterative self-improvement to refine scenarios. Petrovic et al. [13] similarly incorporates LLMs into an autonomous vehicle testing pipeline. Their approach provides the LLM with a formal environment model (metamodel of roads, vehicles, pedestrians, etc.) and standardized requirements as context. The LLM is prompted to produce a concrete test scenario (in a JSON format executable in the CARLA simulator) that satisfies the given requirements. They use the LLM to translate natural-language requirements into Object Constraint Language (OCL) rules—formalizes expected environmental and safety properties. The OCL properties are then checked against the generated test scenario and if required, the feedback is sent to the LLM for correction before the execution of the test scenario. Besides automotive, other related works are emerging in robotics. For example, Wang et al. [22] show that GPT-4 can automatically generate robotic simulation tasks (including environment configurations and goals). They mainly address test scenario generation (test environments and test inputs) rather than directly inferring formal properties or invariants from system specifications. They show that LLMs can handle the environmental context of CPS testing when guided by domain models.
In the broader software systems context, LLMs have been utilized for automated test case generation from various sources of specification. Many of the approaches target conventional software systems (without either ML or any physical components) and have shown promising results in automating unit test creation. Kang et al. [9] present LIBRO, a framework that uses an LLM to generate JUnit tests from bug reports. The goal is to reproduce reported defects automatically as the conventional test generators generally struggle with understanding the semantic intent of a bug report. LIBRO’s performance evaluation on the Defects4J benchmark found that it can produce failing tests for about 33% of bugs and demonstrates that an LLM can interpret natural-language bug descriptions and translate them into fault-revealing code. Another set of work explores using LLMs to generate tests from requirement documents or user stories. Rahman and Zhu [14], leverage GPT-4 to produce test-case specifications (in JSON) directly from high-level requirements and intend to bridge the gap between specifications and executable tests. Some approaches also utilize LLMs within an interactive test generation process. Chen et al. [2] introduce ChatUniTest, an LLM-based unit test generation framework. In their approach, the LLM (Code Llama) drafts a Java unit test; the test is executed to see if it passes or if it exercises the intended code; any errors or unsatisfied goals are fed back for the LLM to repair and refine the test.
Alshahwan et al. [1], report using LLMs to extend and improve existing test suites in an industrial setup—focuses on corner-case inputs that developers missed. Their tool generates additional unit tests to increase coverage of tricky edge conditions. Overall, surveys of the field, e.g., Wang et al., 2024 [21] conclude that LLMs show strong potential in automated testing by reducing the manual effort to write test cases—mainly in code-centric contexts such as unit testing.
Regarding Property-based testing with LLMs, applying LLMs to the generation of PBTs has recently emerged. The most relevant work is by Vikram et al. (2024), which investigates if LLMs write good PBTs [20]. They investigate using GPT-4 and other models to automatically generate PBT code (using the Hypothesis framework in Python) from API documentation. In their setup, the LLM is given the documentation of a library function in natural-language and prompted to generate a property-based test. The generated test produces appropriate random inputs and asserts the documented properties on the outputs. They evaluate the validity (the test must run without errors), soundness (the test assertions should hold for correct implementations and fail for buggy ones) and property coverage (how many distinct expected properties are captured by the test) of the tests.
CPS Challenges and our contribution: The works mentioned above establish a foundation for LLM-driven property-based test generation. In prior studies like Vikram et al.’s [20], the system under test is a software API with no external physical connected components and the LLM did not need to reason about sensors, actuators, or continuous dynamics. But in a cyber-physical system, properties often relate to the interaction between software and the physical components, which are more complex to formalize and test. Environmental mocking becomes a necessity—a model or simulation of the physical environment is required to represent the real world. Recent CPS testing approaches with LLMs (e.g. for autonomous driving [13]) addressed this by restricting the LLM to consider a domain metamodel and produce output in a structured format for a simulator. This helps ensure some basic physical realism in generated scenarios, but it does not guarantee that all relevant properties can be identified or verified. Our approach supports testing and also runtime property-based monitoring. This means the LLM is used to derive property assertions that can also run alongside the deployed system, to check for violations in the runtime. Our approach extends the frontier by applying LLM-driven property-based test generation to CPS, in which both the inference of the generated properties from code documentation and the execution of the corresponding tests must account for the intended CPS programs under test.
6 Conclusion
In this paper, we propose a novel approach for automatically guardrailing cyber-physical system. This approach employs LLMs to generate property-based tests for CPS programs that can be used to monitor CPSs behavior and detect unsafe states. We implement a prototype of this approach in ChekProp and evaluate it on real-world and commonly studied CPSs. We find that ChekProp is applicable on real-world CPSs. More specifically, ChekProp extracts relevant properties, comparable to manually crafted properties, and then generates executable and effective property-based tests that verify these properties.
Our experiments reveal two major challenges for LLM-enabled property-based test generation for CPSs and suggest potential solutions to these challenges. First, given the limited number of public CPS projects, LLMs are not trained on a vast dataset of CPS source code. Consequently, LLMs generated tests may not correctly capture the relation between the API of CPS program and the physical status of the system. For example, while the LLM might recognize that verifying a specific property requires activating the room’s heater, it may not identify the correct method in the CPS code to effect this change. To address this issue, the prompts should contain relevant documents and source code taken from all parts of the project under test. This helps the LLM better understand the API of the CPS code and make the correct use of it.
The second main challenge that we observe is the complexities involved in mocking CPSs. In cyber-physical systems, we deal with the cyber part, the physical part, and the environment. PBTs should run at different times and test the system under various environmental conditions. This requires mocking inputs sometimes from the environment and sometimes from the physical system, and sometimes both. Such mocking demands a correct understanding of the specific environmental inputs to the CPS, their interrelations, and the interface between the cyber and physical components. Our experiments show that LLMs often struggle to capture such an understanding, leading to incorrect generation of PBTs. To address this issue, we suggest including extensive examples of proper mocking scenarios for the CPS under test in the prompt. Prior studies show that such examples can significantly improve the ability of LLMs to generate correct tests [9].
In summary, our study is a solid first step in creating guardrails for CPSs with LLM-generated property-based tests and proves the promising future of this research path.
Acknowledgment
This work is supported by the Trusted Smart Systems pre-study grant at Mälardalen University, as well as Sweden’s Innovation Agency Vinnova through the INTelligent sEcuRity SoluTIons for Connected vEhicles (INTERSTICE) project with grant number 2024-00661 and the Flexible and Secure Modular Automation (FLEXATION) project with grant number 2024-01731.
Appendix A: Example Prompts and Generated PBTs
Here we present two examples of the prompts that ChekProp sends to the LLM and a selected subset of PBTs that are generated based on these prompts.
Listing 5: A PBT generated by ChekProp for the line following robot program.
1@given(
2 left_sensor_value=st.integers(min_value=0, max_value=1),
3 right_sensor_value=st.integers(min_value=0, max_value=1),
4 speed=st.floats(min_value=0.1, max_value=1.0)
5)
6def test_motor_control_based_on_sensor_values(left_sensor_value, right_sensor_value, speed):
7 Device.pin_factory = MockFactory()
8 with LineFollowingRobot(Motor(2, 3, enable=4, pwm=False), Motor(5, 6, enable=7, pwm=False), ...) as lfr:
9 if left_sensor_value > 0:
10 lfr.left_sensor.pin.drive_high()
11 if right_sensor_value > 0:
12 lfr.right_sensor.pin.drive_high()
13 time.sleep(0.1) # Simulate slight delay to trigger motor_speed updates
14
15 if left_sensor_value == 0 and right_sensor_value == 0:
16 assert lfr.right_motor.value == speed
17 assert lfr.left_motor.value == speed
18 elif left_sensor_value == 0 and right_sensor_value == 1:
19 assert lfr.left_motor.value == -speed
20 elif left_sensor_value == 1 and right_sensor_value == 0:
21 assert lfr.right_motor.value == -speed
22 else:
23 assert lfr.left_motor.value != 0 or lfr.right_motor.value != 0
Listing 6: Two PBTs generated by ChekProp for the InputDevice class.
1@given(st.booleans())
2def test_close_releases_pin(pull_up):
3 Device.pin_factory = MockFactory(pin_class=MockPin)
4 device = InputDevice(4, pull_up=pull_up)
5 device.close()
6 assert device.pin is None
7
8@given(st.booleans())
9def test_active_state_affects_is_active_basic(pull_up):
10 Device.pin_factory = MockFactory(pin_class=MockPin)
11 with InputDevice(4, pull_up=pull_up) as device:
12 device.pin.drive_high()
13 assert device.is_active == (not pull_up)
14 device.pin.drive_low()
15 assert device.is_active == pull_up
Figure 4: The prompt used by ChekProp to generate PBTs for the line following robot program.
Figure 5: The prompt used by ChekProp to generate PBTs for the InputDevice class.